in double slit expt the maximum indensity of light would be half (or)twice(or) 4times (or) same
 

Grade:12th pass

1 Answers

Shreyansh Chanani
7 years ago
4 times. Because The formula of maximum intensity is I(max) = (√i+√i)^2. Expanding we get I(max)= i +i + 2i = 4i.